In equilibrium a consumer was buying 5 units of good A and some of good B. His income was Rs 100 and the prices were PA = Rs 8 and PB = Rs 5. The price of good A falls to Rs 5. By how much does his income need to be compensated so that he is able to buy the (old) bundle at the original equilibrium?


1
Expert's answer
2021-07-05T17:40:22-0400

According to the given information, units of B consumed by the consumer would be:

"M=PA(QA)+PB(QB)\\

100=8(5)+5(QB)\\

100-40=5(QB)\\

\frac{60}{5}=QB\\

12=QB"

Therefore, original bundle consumed by the consumer would be (5,12)

Here, it is given that the price of good A falls to Rs 5 and consumer wants to consume the same bundle as earlier. Therefore, using the given information, new income to consume original bundle would be:

"M=PA(QA)+PB(QB)\\

M=5(5)+5(12)\\

M=25+60\\

M=85"

Therefore, he needs to compensate Rs 15 (Rs100 - Rs85) from his income so that he is able to buy the (old) bundle at the original equilibrium.